OpenStudy (anonymous):

help solve the given equation sin^2x-2sin x-3=0

OpenStudy (anonymous):

you want to? simplify? find x? what? ^_^

OpenStudy (dumbcow):

use substitution assign u = sinx then u^2 -2u -3 = 0 Factor -3*1 = -3 -3+1 = -2 (u-3)(u+1) =0 u=3 u=-1 replace u back with sinx and solve for x sinx = 3, 3 is outside range of function so no solution sinx = -1 look at unit circle sinx is -1 when x =3pi/2